# SynTQA: Synergistic Table-based Question Answering via Mixture of Text-to-SQL and E2E TQA

### v1

@@ 1 Introduction
Table QA (TQA) takes a question and a table, and finds an answer based on the evidence from the table Pasupat and Liang (2015). With the help of large scale datasets Zhong et al. (2017); Yu et al. (2018, 2019); Shi et al. (2020a), state-of-the-art (SOTA) TQA systems primarily focus on two approaches: semantic parsing (Text-to-SQL) that predicts a SQL query as intermediate semantic representation of the question, and then executes the SQL to find the answer (Wang et al., 2020; Scholak et al., 2021; Li et al., 2023a); end-to-end system (E2E TQA) that directly generates the answer from models pre-trained on table corpora, imitating human-like reasoning on questions and tables (Pasupat and Liang, 2015; Iyyer et al., 2017; Gupta et al., 2023). Despite serving for a similar purpose, it’s unclear what advantages these approaches have and their potential synergy.
@@
To answer these questions, we first (re)evaluate SOTA Text-to-SQL models, i.e., T5 Raffel et al. (2020), GPT (OpenAI, 2023), and DIN-SQL (Pourreza and Rafiei, 2023a), as well as E2E TQA models, i.e., TaPEx (Liu et al., 2022), OmniTab (Jiang et al., 2022), and GPT, on benchmark datasets WTQ (Pasupat and Liang, 2015) and WikiSQL (Zhong et al., 2017). The experiments show that while both Text-to-SQL and E2E TQA approaches are adept at simple questions, they have complementary strengths for complex questions and tables, as shown in Figure 1 (top): Text-to-SQL is more proficient in numerical reasoning and processing long tables while E2E TQA is better at ambiguous questions, complex schema and contents.
@@ 2 Table Question Answering Task
Table QA problem is originally framed as semantic parsing, also known as Text-to-SQL parsers, where a parser takes both question and table header as input, and predicts a SQL query that is directly executable to get the answer. Early neural sequence-to-sequence parsers (Guo et al., 2019; Wang et al., 2020; Rubin and Berant, 2021) encode question/schema with attention mechanism and uses SQL grammar to guide the decoding process. Recent approaches take advantages of pre-trained models, and they either fine-tune Wang et al. (2018); Scholak et al. (2021) or prompt Gao et al. (2023); Pourreza and Rafiei (2023b) large models for Text-to-SQL parsing.
@@ 3.1 Experimental Setup
WTQ comprises 22,033 instances with a diverse array of intricate questions and tables. Squall (Shi et al., 2020a) annotates 11,276 WTQ instances with pre-processed tables and SQL queries. [^1] We train Text-to-SQL models on Squall and test on WTQ, as $20\%$ of WTQ questions lack SQL annotations and cannot be answered by Text-to-SQL. Compared with classic datasets, e.g., WikiSQL and Spider (Yu et al., 2018), designed for SQL prediction on well-maintained databases, WTQ contains complex tables and questions which are difficult to answer with SQL queries. As a large portion of Spider tables does not have table content, we use WikiSQL to validate the generalizability of our findings, which contains 80,654 instances.
@@
We evaluate SOTA models that have publicly available source code or APIs: Text-to-SQL includes T5 (Raffel et al., 2020), GPT (OpenAI, 2023), and DIN-SQL (Pourreza and Rafiei, 2023a); E2E TQA includes TaPEx (Liu et al., 2022), OmniTab Jiang et al. (2022), and GPT (OpenAI, 2023). As Text-to-SQL models often generate invalid SQL queries (Lin et al., 2020; Scholak et al., 2021), we devise a post-processing module to screen table content, rectify query misspellings, identify the closest string values, and resolve mismatches. For fine-tuned models, we choose the large version. For fair comparison, we report answer string exact match (EM) accuracy.
@@ 3.2 Results
According to Table 1, prompting methods (i.e., GPT and DIN-SQL) underperform fine-tuned models in table understanding on WTQ and WikiSQL, aligning with findings in (Li et al., 2023b; Liu et al., 2023). Thus, we primarily focus on fine-tuned Text-to-SQL and E2E TQA models. Best Text-to-SQL and E2E TQA models achieve comparable accuracy, but notably, 27.6% of WTQ and 11.7% of WikiSQL questions were correctly answered exclusively by either Text-to-SQL or E2E TQA. It implies that models excel in tackling different types of table-based questions. To further investigate the strengths and weaknesses, we analyze 200 erroneous cases summarized in Figure 2 (see detailed breakdown in Appendix B).
@@
It is evident in Figure 2 (A) that 61% of E2E TQA error cases involve arithmetic operations including counting, summation, averaging, and subtraction. Despite existing E2E TQA approaches (Herzig et al., 2020; Eisenschlos et al., 2020) have incorporated a separate aggregation operator into model design, the range of supported operations is limited with suboptimal performance. In contrast, Text-to-SQL provides more accurate and consistent results for arithmetic operations through symbolic reasoning (Cheng et al., 2023; Liu et al., 2023).
@@
As pointed out by (Shi et al., 2020b), there are cases where SQL queries are insufficiently expressive. According to Figure 2 (E), Text-to-SQL cannot answer questions related to phrases “approximate”, while E2E TQA is about to find the answers.
@@ 4.2 Results
According to Table 1 (Bottom), our ensemble models exhibit substantial improvement over individual models. They achieve comparable performance with recent tool-based LLMs on WTQ while saving computational costs, e.g., Dater (Ye et al., 2023) 65.9% and Mix SC (Liu et al., 2023) 73.6%. As our findings are orthogonal to these methods, we demonstrate a case integrating the concept of Mix SC in Appendix G. The effectiveness can be attributed to selectors’ high success rate (nearly 80%) in selecting correct answers. Notably, the confidence of Text-to-SQL and E2E TQA models is the most impactful feature for SynTQA (RF).
@@ 4.3 SQL Annotation Efficiency
Since manually creating SQL annotations can be costly (Shi et al., 2020a), we conducted experiments to study how the accuracy improvement varies with different amounts of SQL annotations, using the feature-based selector in the WTQ dataset. The answers are assumed to be always fully available, leading to a stable performance of E2E TQA.
@@ 5 Other Related Work
With LLMs’ strong textual reasoning and tool-use capabilities, recent Table QA methods (Cheng et al., 2023; Ye et al., 2023; Liu et al., 2023) call executable programs (e.g., SQL and Python) as needed to retrieve relevant contexts, facilitating reasoning. We provide an alternative ensemble approach that does not rely on computationally expensive LLMs.
@@ 6 Conclusion
Although OmniTab is pre-trained for E2E TQA, T5 is not a model specifically designed for Text-to-SQL. Most Text-to-SQL models are tailored for the Spider dataset (Wang et al., 2018; Rubin and Berant, 2021; Scholak et al., 2021). Table or passage retrievers (Karpukhin et al., 2020; Herzig et al., 2021) can be applied to select certain rows and columns before truncating the long tables which might improve E2E TQA performance. As for SynTQA (GPT), we constrain GPT to select an answer from candidates, which abandons its capability to provide a different answer when both candidates are wrong. In more challenging datasets which necessitate both textual and tabular data (Chen et al., 2020; Zhu et al., 2021), our method may not be as flexible and effective as tool-based LLMs (Li et al., 2023c; Asai et al., 2023).
@@
SynTQA were developed using WTQ (Pasupat and Liang, 2015), Squall (Shi et al., 2020a), WikiSQL (Zhong et al., 2017), and RobuT Zhao et al. (2023), which are publicly available under the licenses of CC BY-SA 4.0444https://creativecommons.org/licenses/by-sa/4.0/, BSD 3-Clause [^5] https://opensource.org/license/bsd-3-clause, and MIT [^6] https://opensource.org/license/mit. We used 4 NVIDIA Quadro RTX8000 GPUs to fine-tune models. SynTQA (RF) and SynTQA (GPT) were constructed and executed solely using CPU. SynTQA (GPT) relies on OpenAI API and using other GPT versions will lead to varied performance. No manual annotation and human study are involved in this study.
@@ Acknowledgements
This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix D LLM-based Table QA Models
This section presents the evaluation of LLM-based E2E TQA and Text-to-SQL models. To optimize the cost, we use gpt-3.5-turbo-0125 for all models including E2E TQA, Text-to-SQL, and SynTQA (GPT) selector. For LLM-based E2E TQA, we follow the direct prompting (zero-shot) approach implemented by (Liu et al., 2023). For LLM-based Text-to-SQL, we incorporate 8 examples from dev set in the prompting to showcase the target style of SQL queries.
@@ F.3 Alignment Module
F.4 Contradiction Module
@@ Appendix G Integrating Self-Consistency
Following (Liu et al., 2023), we incorporate the Self-Consistency in our Text-to-SQL and E2E TQA base models. To generate 5 candidate answers for each model, we perturb the input table schema for the Text-to-SQL model and conduct top-k sampling ($k=50$) for the E2E TQA model. Among five candidates, we choose one following the rule of maximum voting. Lastly, our RF classifier determined the final answer based on designed features.

### v2

@@ 1 Introduction
Table QA (TQA) takes a question and a table, and finds an answer based on the evidence from the table Pasupat and Liang (2015). With the help of large scale datasets Zhong et al. (2017); Yu et al. (2018, 2019); Shi et al. (2020), state-of-the-art (SOTA) TQA systems primarily focus on two approaches: semantic parsing (Text-to-SQL) that predicts a SQL query as intermediate semantic representation of the question, and then executes the SQL to find the answer (Wang et al., 2020; Scholak et al., 2021; Li et al., 2023a); end-to-end system (E2E TQA) that directly generates the answer from models pre-trained on table corpora, imitating human-like reasoning on questions and tables (Pasupat and Liang, 2015; Iyyer et al., 2017; Gupta et al., 2023). Despite serving for a similar purpose, it’s unclear what advantages these approaches have and their potential synergy.
@@
To answer these questions, we first (re)evaluate SOTA Text-to-SQL models, i.e., T5 Raffel et al. (2020), GPT (OpenAI, 2023), and DIN-SQL (Pourreza and Rafiei, 2024), as well as E2E TQA models, i.e., TaPEx (Liu et al., 2022), OmniTab (Jiang et al., 2022), and GPT, on benchmark datasets WTQ (Pasupat and Liang, 2015) and WikiSQL (Zhong et al., 2017). The experiments show that while both Text-to-SQL and E2E TQA approaches are adept at simple questions, they have complementary strengths for complex questions and tables, as shown in Figure 1 (top): Text-to-SQL is more proficient in numerical reasoning and processing long tables while E2E TQA is better at ambiguous questions, complex schema and contents.
@@ 2 Table Question Answering Task
Table QA problem is originally framed as semantic parsing, also known as Text-to-SQL parsers, where a parser takes both question and table header as input, and predicts a SQL query that is directly executable to get the answer. Early neural sequence-to-sequence parsers (Guo et al., 2019; Wang et al., 2020; Rubin and Berant, 2021) encode question/schema with attention mechanism and uses SQL grammar to guide the decoding process. Recent approaches take advantages of pre-trained models, and they either fine-tune Wang et al. (2018); Scholak et al. (2021) or prompt Gao et al. (2024); Pourreza and Rafiei (2024) large models for Text-to-SQL parsing.
@@ 3.1 Experimental Setup
WTQ comprises 22,033 instances with a diverse array of intricate questions and tables. Squall (Shi et al., 2020) annotates 11,276 WTQ instances with pre-processed tables and SQL queries. [^1] We train Text-to-SQL models on Squall and test on WTQ, as $20\%$ of WTQ questions lack SQL annotations and cannot be answered by Text-to-SQL. Compared with classic datasets, e.g., WikiSQL and Spider (Yu et al., 2018), designed for SQL prediction on well-maintained databases, WTQ contains complex tables and questions which are difficult to answer with SQL queries. As a large portion of Spider tables does not have table content, we use WikiSQL to validate the generalizability of our findings, which contains 80,654 instances.
@@
We evaluate SOTA models that have publicly available source code or APIs: Text-to-SQL includes T5 (Raffel et al., 2020), GPT (OpenAI, 2023), and DIN-SQL (Pourreza and Rafiei, 2024); E2E TQA includes TaPEx (Liu et al., 2022), OmniTab Jiang et al. (2022), and GPT (OpenAI, 2023). As Text-to-SQL models often generate invalid SQL queries (Lin et al., 2020; Scholak et al., 2021), we devise a post-processing module to screen table content, rectify query misspellings, identify the closest string values, and resolve mismatches. For fine-tuned models, we choose the large version. For fair comparison, we report answer string exact match (EM) accuracy.
@@ 3.2 Results
According to Table 1, prompting methods (i.e., GPT and DIN-SQL) underperform fine-tuned models in table understanding on WTQ and WikiSQL, aligning with findings in (Li et al., 2024; Liu et al., 2024). Thus, we primarily focus on fine-tuned Text-to-SQL and E2E TQA models. Best Text-to-SQL and E2E TQA models achieve comparable accuracy, but notably, 27.6% of WTQ and 11.7% of WikiSQL questions were correctly answered exclusively by either Text-to-SQL or E2E TQA. It implies that models excel in tackling different types of table-based questions. To further investigate the strengths and weaknesses, we analyze 200 erroneous cases summarized in Figure 2 (see detailed breakdown in Appendix B).
@@
It is evident in Figure 2 (A) that 61% of E2E TQA error cases involve arithmetic operations including counting, summation, averaging, and subtraction. Despite existing E2E TQA approaches (Herzig et al., 2020; Eisenschlos et al., 2020) have incorporated a separate aggregation operator into model design, the range of supported operations is limited with suboptimal performance. In contrast, Text-to-SQL provides more accurate and consistent results for arithmetic operations through symbolic reasoning (Cheng et al., 2023; Liu et al., 2024).
@@
As pointed out by (Shi et al., 2020), there are cases where SQL queries are insufficiently expressive. According to Figure 2 (E), Text-to-SQL cannot answer questions related to phrases “approximate”, while E2E TQA is about to find the answers.
@@ 4.2 Results
According to Table 1 (Bottom), our ensemble models exhibit substantial improvement over individual models. They achieve comparable performance with recent tool-based LLMs on WTQ while saving computational costs, e.g., Dater (Ye et al., 2023) 65.9% and Mix SC (Liu et al., 2024) 73.6%. As our findings are orthogonal to these methods, we demonstrate a case integrating the concept of Mix SC in Appendix G. The effectiveness can be attributed to selectors’ high success rate (nearly 80%) in selecting correct answers. Notably, the confidence of Text-to-SQL and E2E TQA models is the most impactful feature for SynTQA (RF).
@@ 4.3 SQL Annotation Efficiency
Since manually creating SQL annotations can be costly (Shi et al., 2020), we conducted experiments to study how the accuracy improvement varies with different amounts of SQL annotations, using the feature-based selector in the WTQ dataset. The answers are assumed to be always fully available, leading to a stable performance of E2E TQA.
@@ 5 Other Related Work
With LLMs’ strong textual reasoning and tool-use capabilities, recent Table QA methods (Cheng et al., 2023; Ye et al., 2023; Liu et al., 2024) call executable programs (e.g., SQL and Python) as needed to retrieve relevant contexts, facilitating reasoning. We provide an alternative ensemble approach that does not rely on computationally expensive LLMs.
@@ 6 Conclusion
Although OmniTab is pre-trained for E2E TQA, T5 is not a model specifically designed for Text-to-SQL. Most Text-to-SQL models are tailored for the Spider dataset (Wang et al., 2018; Rubin and Berant, 2021; Scholak et al., 2021). Table or passage retrievers (Karpukhin et al., 2020; Herzig et al., 2021) can be applied to select certain rows and columns before truncating the long tables which might improve E2E TQA performance. As for SynTQA (GPT), we constrain GPT to select an answer from candidates, which abandons its capability to provide a different answer when both candidates are wrong. In more challenging datasets which necessitate both textual and tabular data (Chen et al., 2020; Zhu et al., 2021), our method may not be as flexible and effective as tool-based LLMs (Li et al., 2023b; Asai et al., 2024).
@@
SynTQA were developed using WTQ (Pasupat and Liang, 2015), Squall (Shi et al., 2020), WikiSQL (Zhong et al., 2017), and RobuT Zhao et al. (2023), which are publicly available under the licenses of CC BY-SA 4.0444https://creativecommons.org/licenses/by-sa/4.0/, BSD 3-Clause [^5] https://opensource.org/license/bsd-3-clause, and MIT [^6] https://opensource.org/license/mit. We used 4 NVIDIA Quadro RTX8000 GPUs to fine-tune models. SynTQA (RF) and SynTQA (GPT) were constructed and executed solely using CPU. SynTQA (GPT) relies on OpenAI API and using other GPT versions will lead to varied performance. No manual annotation and human study are involved in this study.
@@ Acknowledgements
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore. Siyue Zhang and Chen Zhao were supported by Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning, NYU Shanghai. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.
@@ Appendix D LLM-based Table QA Models
This section presents the evaluation of LLM-based E2E TQA and Text-to-SQL models. To optimize the cost, we use gpt-3.5-turbo-0125 for all models including E2E TQA, Text-to-SQL, and SynTQA (GPT) selector. For LLM-based E2E TQA, we follow the direct prompting (zero-shot) approach implemented by (Liu et al., 2024). For LLM-based Text-to-SQL, we incorporate 8 examples from dev set in the prompting to showcase the target style of SQL queries.
@@ F.3 Alignment Module
F.4 Comparison Module
F.5 Contradiction Module
@@ Appendix G Integrating Self-Consistency
Following (Liu et al., 2024), we incorporate the Self-Consistency in our Text-to-SQL and E2E TQA base models. To generate 5 candidate answers for each model, we perturb the input table schema for the Text-to-SQL model and conduct top-k sampling ($k=50$) for the E2E TQA model. Among five candidates, we choose one following the rule of maximum voting. Lastly, our RF classifier determined the final answer based on designed features.